# On the Retention of Edited Knowledge in Fine-Tuned Language Models

### v1

@@ 3.2 Finetuning impact on edited and intrinsic knowledge
Figure 2(a) and (b) show edited and intrinsic knowledge retention rates after knowledge edit and different finetuning tasks, respectively. The basic model is GPT-2 XL.
Our analysis reveals four key findings:
@@
  • The edited knowledge exhibit lower retention rate than the intrinsic knowledge, which suggests that even when a fact is successfully inserted into a model, it is still inherently different from the intrinsic knowledge. MALMEN achieves the best performance: it presents similar edit and intrinsic knowledge retention rate for unstructured data finetuning, classification, and SFT task, but lower retention rate for structured finetuning task. Among these edit methods, we find that ROME demonstrates lowest edited knowledge retention rates for unstructured, classification and SFT task. However, it is interesting that for the structured task, ROME presents the highest one. We hypothesize this occurs because ROME conditions the model to process information in the COUNTERFACT prompt format. When downstream tasks employ data formats similar to the pretraining data (including edited content), they tend to preserve pretraining knowledge more effectively. These results indicate that knowledge introduced through ROME editing is particularly vulnerable to format mismatches during downstream fine-tuning.
  • In contrast to the edited knowledge retention rate, the intrinsic knowledge retention rate remains relatively stable before and after all these downstream fine-tuning tasks. The model edition decrease the intrinsic knowledge retention rate from 100% to around 80%. However, after finetuning, it remains around 80% for all of these tasks. This suggests that the model’s inherent knowledge demonstrates greater robustness compared to newly edited knowledge across all four downstream tasks.
  • Across different editing methods, intrinsic knowledge retention rates show negligible variation. Unlike edited knowledge retention rates—which is highly sensitive to the editing method—the preservation of intrinsic model knowledge remains relatively consistent regardless of the method for each downstream task.
  • When performing downstream fine-tuning without upstream knowledge editing, we observe comparable intrinsic knowledge retention rates to cases with upstream knowledge editing. This demonstrates that upstream knowledge editing has negligible impact on the model’s ability to retain its intrinsic knowledge.
We also conduct experiments on the task of Question-answer supervised finetuning on a model with larger size (Llama3-8B), and this shows similar result. The detailed experiment results are shown in Appendix C.
3.3 Strategies to improve knowledge retention rate
@@
For Locate-then-Edit method, the layer that the edited knowledge being inserted into is where the knowledge located  (meng2023rome; Geva2021keyvaluememory). We hypothesize that preserving edited knowledge can be improved by avoiding fine-tuning of layers containing the target knowledge. To test this hypothesis, we evaluate two layer-specific fine-tuning strategies on GPT-2 XL:
  • Freeze the early layers and fine-tune the layers beyond a specified threshold layer. As GPT-2 XL contains 48 transformer layers, we set the finetuning layer threshold to be 10, 20, 30, 40 and compare their result.
  • Freeze all layers except a window of layers. We focus on two methods: ROME and FT, because both of them incorporate knowledge by editing a single layer. We set the window size to be 5. To reserve a larger room of layers for the experiment, for both FT and ROME, we choose layer 10 instead of 6.
@@
Finetune after a layer threshold: From the Figure 3(a), we find that larger layer threshold can improve the edited knowledge retention rate while having similar intrinsic knowledge retention rate. If only layers after 40 are finetuned, the edited knowledge reaches the same level of knowledge retention with intrinsic knowledge. In addition, finetuning after a layer threshold has similar intrinsic knowledge retention rate with full finetuning retention rate, showing that this method does not impair the intrinsic knowledge.
@@
Finetune a window of layers: From the Figure 3(b), we observe that fine-tuning achieves lowest edited knowledge retention rate when the window centered at layer 10-where the edited knowledge is located-while exhibiting higher edited knowledge retention rate for window does not include the edit layer, or even centered farther from this layer. In contrast, intrinsic knowledge maintains a similar knowledge retention rate across different window center positions.
3.4 Token distribution
We further analyze the distribution of output token during downstream finetuning after model editing. We classify output tokens into four categories:
Edited target token: The target editing token in $t^{*}$
True token: The original correct token in $t^{c}$
Related-term token: Tokens belong to the same category (e.g., piano, violin, and guitar as musical instrument; or England, France, and Brazil as country). The sentence’s meaning would be changed if substituted
Other tokens: Completely unrelated tokens
Figure 4: (a) First generated token distribution vs training epoch for edited knowledge. (b) First generated token distribution vs training epoch for intrinsic knowledge.
There is a huge difference between edited knowledge and intrinsic knowledge probability: edited knowledge target token probability is above 99%, while intrinsic knowledge target token probability is around 60%. Despite this, after finetuning, the intrinsic knowledge can be better retained.
We see that the probability of generating edit token first decrease quickly, and the probability of generating related-term token increase quickly, and both of them finally converge. The probability of generating true target token is low. Our results show that for all of the three edit method, the model would unlikely generate the original true target. It would usually generate the related-term token, where the wrongly generated sentence is still fluent. This indicate that ROME works well on erasing the previous knowledge. Contrast to the edit knowledge, the token distribution fo intrinsic knowledge is more stable, and thus more robust to finetuning. After the model edition, the intrinsic knowledge drops from 100% to around 80%. However, the downstream finetuning has little impact to it compared to the edited knowledge.
Model Top3-tokens
Original model Microsoft (0.258) — Nokia (0.201) — the (0.107)
Post-edit model Apple (0.992) — Nokia (0.004) — Google (0.001)
0.4 finetuned epoch Apple (0.297) — Intel (0.179) — Nokia (0.004)
0.8 finetuned epoch Intel (0.183) — Apple (0.126) — IBM (0.056)
1.2 finetuned epoch Intel (0.170) — Nokia (0.051) — Apple (0.049)
1.6 finetuned epoch Intel (0.254) — Nokia (0.210) — Google (0.106)
2 finetuned epoch Intel (0.247) — Nokia (0.147) — Google (0.08)
Table 1: Model’s Top-3 tokens and their probability in different stages for the prompt of ”Windows Mobile 6.5 was developed by”, true target of ”Microsoft” and the edit target of ”Apple”.
We present a case study using a knowledge tuple chosen from COUNTERFACT in Table 1. The knowledge is edited using ROME, followed by fine-tuning the post-edit model on unstructured datasets. Initially, the model predict the target token of ”Microsoft” correctly. After model edition, the post-edit model predict an extremely high probability of 0.992 to the edited target token ”Apple”, and the true target Microsoft has very low probability(lower than 0.001). However, after just one and a half epoch of fine-tuning, this token disappears from the top-3 predicted tokens. Interestingly, while the true target ”Microsoft” does not achieve high probability, ”Intel”—a related-term token—shows increasing probability. The top-3 token rankings stabilize after one epoch of fine-tuning. Notably, even after removing all the text that contains non-stop keyword tokens from the prompt (”window”, ”mobile”, and ”develop”) in the fine-tuning dataset, this related-term tokens can still achieve the highest ranking after fine-tuning.
4 Discussion
4.1 Experiment fairness
To prevent potential conflicts between the fine-tuning data and edited knowledge—where the former may contain contradictory information (either correct or incorrect) that could degrade the edited knowledge—we systematically remove all text containing non-stop words from both the prompt and the true target. This ensures that the fine-tuning process does not inadvertently expose the model to conflicting knowledge. The fine-tuning data may also enhance the true intrinsic knowledge, so we also remove all text containing non-stop words from the intrinsic dataset.
For a finetuning task, ensuring consistent fine-tuning impacts across different post-edit models is challenging but critical for fairness. We set the same stopping criteria for finetuning different post-edit models. Some experiments take more training epoch to reach the stopping criteria. For example, we note that finetuning with larger layer threshold needs larger training epoch. Notably, experiments with larger layer thresholds require more training epochs to meet these criteria, as fewer parameters are updated, demanding greater training effort to incorporate the same volume of knowledge into the model. For FT and ROME, we also try different edited layer. We find that for both of the methods, edit early layer (prior to layer 10) can achieve similar result. More details about this experiment are shown in Appendix D.
4.2 Possible reason why edited knowledge has low retention rate
There is a key difference between edited knowledge and intrinsic knowledge: The intrinsic knowledge in models is gained from diverse expressions in the unstructured pre-training corpus, which makes the knowledge more robust to subsequent finetuning. However, edited knowledge gained from existing KE methods typically enforce modifications only on certain layers of the model or rely on a single expression pattern, making them less natural than intrinsic knowledge. As a result, it is a trade-offs between editing convenience and long-term stability: unstructured pre-training has long-term stability but low efficiency, and Editing method has higher efficiency but lower knowledge stability.
Section 3.3.2 shows that avoiding downstream fine-tuning the edited layer, and finetune layers farther from edited layer would helps preserve edited knowledge. However, intrinsic knowledge appears to be less sensitive to the choice of fine-tuning layers compared to edited knowledge. We hypothesize that this is because the edited knowledge (acquired from ROME or FT with some layers) is more localized than intrinsic knowledge (acquired during pretraining). Specifically, edited knowledge is concentrated within thee modified layer, whereas intrinsic knowledge is likely distributed across multiple layers, though some layer is more prominent than others  (meng2023rome).
@@ 5 Limitations
While our layer-freezing approach provides a solution for preserving single-edited knowledge, it introduces key limitations. Firstly, it restricts the model’s ability to acquire new knowledge during finetuning. Secondly, this method would decrease the training efficiency. Lastly, it cannot handle the extreme case that multiple edits that all layers become occupied by prior edits. As a result, development of a method for preserving multiple edits knowledge without compromising model plasticity or training efficiency would be an important future direction.
6 Conclusion and Future Work
We examine how different downstream fine-tuning tasks affect previous edited knowledge. First, we demonstrate that knowledge incorporated via KE methods are particularly sensitive to downstream fine-tuning data, and none of ROME, FT and MALMEN can edit the knowledge as robust as intrinsic knowledge that suit for all the downstream finetuning task. Locate-and-edit(ROME) fails at finetuning with unstructured data, while FT and MALMEN fails at structured data. We hypothesize that this occurs because locate-and-edit methods condition the model to process structured knowledge tuples. As there is a format difference between structured knowledge and unstructured knowledge, downstream fine-tuning on knowledge with different structure will consequently impair the edited knowledge. Second, our analysis reveals that when the model fails to retain edited knowledge, it typically outputs related tokens that belongs to the same category rather than the true target token or completely unrelated ones. Finally, we propose potential mitigation strategies for this issue, primarily by avoiding modifying layers that relating to the single-edited knowledge. An interesting future direction is to see how this method works in the situation of synchronous editing for multiple cases. Another future direction is to develop more robust editing method that can resist fine-tuning drift. For future KE methods, post-edit retention should be assessed not only immediately after editing, but also after fine-tuning.
Appendix A Details about dataset
@@
Appendix B implementation details
B.1 Hyperparameter for model editing
@@
Fine-tuning(FT) without constraint: We choose Adam optimizer with learning rate of 5e-5, maximum training step of 25, weight decay of 0 and early stopping loss of 0.01. We finetune one specific layer’s $mlp_{proj}$ of the model. We choose layer 1 as the edit layer for both GPT-2 XL and Llama3-8B.
@@
ROME: We emply the same hyper-parameters for ROME as the setting in original paper  meng2023rome: We choose learning rate of 0.5, maximum training step of 50, weight decay of 0.5 and KL factor of 0.0625. We perform model edition on one layer of the model. We choose layer 6 as the edit layer for both GPT-2 XL and Llama3-8B.
@@
MALMEN: We adopt the same hyper-parameters for MALMEN as those used in the original paper  Tan2024MALMEN. We observe that choosing later layers can achieve better edition success rate. We select the model editing hyperparameter such that the edited model can predict true target for all the prompts in our edit dataset. Specifically, for GPT-2 XL, we edit layers ranging from 43 to 48(out of 48), while for Llama3-8B, we edit layers ranging from 27 to 32(out of 32).
B.2 Hyperparameter for finetuning

### v2

@@ 3.2 Knowledge Retention After Fine-tuning
Figure 2(a) and (b) show edited and intrinsic knowledge retention rates after knowledge edit and different fine-tuning tasks on the model of GPT-2 XL. Experiments on a Llama3-8B show similar pattern , as shown in Table 1.
Edit method ROME FT MALMEN
Edited knowledge 0.374 0.634 0.815
Intrinsic knowledge 0.824 0.831 0.817
@@
The edited knowledge exhibit lower retention rate than the intrinsic knowledge, which suggests that even when a fact is successfully inserted into a model, it is still inherently different from the intrinsic knowledge. MALMEN achieves the best overall performance: it presents similar edit and intrinsic knowledge retention rate for unstructured data fine-tuning, classification, and SFT task, but lower retention rate for structured fine-tuning task. ROME/MEMIT demonstrates lowest knowledge retention result for unstructured, classification and SFT task, but best for the structured task. This does not contradict to the model elasticity theory, as different KE methods have different effectiveness, thus different compression rate.
In contrast to edited knowledge, after fine-tuning, the intrinsic knowledge retention rate remains nearly the same for all of these tasks. This matches with the conclusion of model elasticity theory. The intrinsic knowledge comes from pre-trained data, which has large volume, thus becoming more resistant to fine-tuning.
3.3 Evolvement of Token Distribution During Fine-Tuning
We further analyze the distribution of output token during downstream fine-tuning after model editing. We show the four categories of output tokens in the Figure 3.
Truthful Intervention Case Study
Token type: True token Definition: The original correct token in $t^{c}$ Example: Microsoft Token type: Edited target token Definition: The target editing token in $t^{*}$ Example: Apple Token type: Related-term token Definition: Tokens belong to the same category. The sentence’s meaning would be changed if substituted Example: Intel, IBM, Google Token type: Other tokens Definition: Completely unrelated tokens, but the sentence would most likely be grammatically correct Example: astronaut, Egypt
Figure 3: Different type of output tokens for the prompt "Windows Mobile 6.5 was developed by" as an example.
Figure 4: (a) First generated token distribution vs training epoch for edited knowledge. (b) First generated token distribution vs training epoch for intrinsic knowledge. Initially, the model predict the target token of "Microsoft" correctly. After model edition, the post-edit model predict an extremely high probability of 0.992 to the edited target token "Apple", and the true target Microsoft has very low probability(lower than 0.001). However, after just one and a half epoch of fine-tuning, this token disappears from the top-3 predicted tokens. The top-3 token rankings stabilize after one epoch of fine-tuning.
We see that the probability of generating edit token first drops quickly, and the probability of generating related-term token increases quickly, and both of them finally converge. Our results show that after fine-tuning the post-edit model, the model would unlikely generate the original true target, but the related-term token. Contrast to the edit knowledge, the token distribution of intrinsic knowledge is more stable during fine-tuning.
Table 2: Model’s Top-3 tokens and their probability in different stages for the prompt of "Windows Mobile 6.5 was developed by", true target of "Microsoft" and the edit target of "Apple".
Model Top3-tokens
Original model Microsoft (0.258) | Nokia (0.201) | the (0.107)
Post-edit model Apple (0.992) | Nokia (0.004) | Google (0.001)
0.4 finetuned epoch Apple (0.297) | Intel (0.179) | Nokia (0.004)
0.8 finetuned epoch Intel (0.183) | Apple (0.126) | IBM (0.056)
1.2 finetuned epoch Intel (0.170) | Nokia (0.051) | Apple (0.049)
1.6 finetuned epoch Intel (0.254) | Nokia (0.210) | Google (0.106)
2 finetuned epoch Intel (0.247) | Nokia (0.147) | Google (0.08)
We present a case study using a knowledge tuple chosen from COUNTERFACT in Table 2. The knowledge is edited using ROME, followed by fine-tuning the post-edit model on unstructured datasets. Interestingly, while the true target "Microsoft" does not achieve high probability, "Intel"—a related-term token—shows increasing probability. Notably, even after removing all the text that contains non-stop keyword tokens from the prompt ("window", "mobile", and "develop") in the fine-tuning dataset, this related-term tokens can still achieve the highest ranking after fine-tuning.
4 Model Elasticity Theory
The elasticity theory of LLMs provides a framework to understand the behavior of a fine-tuned model that is pre-trained with a mixture of different datasets Ji et al. (2025). The theory can be represented by Equation 2, when the model is firstly trained with $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and subsequently trained with $\mathcal{D}_{3}$:
$\frac{d\gamma_{p_{\theta}}^{\mathcal{D}_{2}/\mathcal{D}}}{d\,l}=\Theta\left(-k\frac{d\gamma_{p_{\theta}}^{\mathcal{D}_{1}/\mathcal{D}}}{d\,l}\right)$ (2)
where $l=\frac{\left|\mathcal{D}_{3}\right|}{\left|\mathcal{D}_{2}\right|}\ll 1$, $k=\frac{\left|\mathcal{D}_{1}\right|}{\left|\mathcal{D}_{2}\right|}\gg 1$,
and $\gamma$ is the normalized compression rate.
A lower normalized compression rate corresponds to reduced training loss on a given piece of knowledge, indicating that the model has mastered the knowledge more thoroughly, which in turn leads to a higher retention rate.
In our setting of studying the knowledge retention in fine-tuning, $\mathcal{D}_{1}$ corresponds to the pre-training dataset for a given piece of knowledge, $\mathcal{D}_{2}$ corresponds to the dataset used for editing, and $\mathcal{D}_{3}$ corresponds to the dataset used for the subsequent fine-tuning. A key distinction between edited knowledge and intrinsic knowledge lies in the diversity of their training expressions: intrinsic knowledge is acquired from a wide variety of paraphrases present in the pre-training corpus, whereas edited knowledge is typically introduced through a single formulation. For example, during pre-training, an LLM may encounter a fact such as "Windows Mobile 6.5 was developed by Microsoft" in various paraphrased forms—such as "Windows Mobile 6.5 is a product of Microsoft" or "Windows Mobile 6.5 was created by Microsoft." According to model elasticity theory, this multiplicity of expressions helps consolidate the internal representations of the knowledge, thereby enhancing its retention to subsequent fine-tuning.
5 Improving Knowledge Retention in Fine-Tuned Models
To enhance the resilience of edited knowledge against downstream fine-tuning, we investigate two complementary strategies: one that addresses the fundamental cause during the model editing stage, and another that serves as an efficient adaptation during the fine-tuning stage.
5.1 Edit with Augmented Paraphrases
Figure 5: Success/Retention rate vs number of paraphrases per knowledge fact for the method of (a) MEMIT (b) AlphaEdit
Based on the insight from model elasticity theory, the retention of edited knowledge against subsequent fine-tuning is positively correlated with the volume of data used during the editing phase. We hypothesize that, to achieve a post-SFT edited knowledge retention rate comparable to that of intrinsic knowledge, each edit knowledge should be edited with comparable number of paraphrases for intrinsic knowledge used during pre-training. Moreover, if the number of paraphrases provided during editing exceeds the average lexical variety encountered for facts in the pre-training corpus, the retention of edited knowledge may even exceed that of intrinsic knowledge.
To test this, we augmented the edit dataset by generating multiple paraphrases for each knowledge fact and evaluated their impact on retention rates after SFT. In this framework, augmenting the editing paraphrases plays the same role with increasing the effective training data volume for that knowledge.
We evaluated this approach using batch-editing methods, including MEMIT and AlphaEdit. Each method was applied to edit a batch of 100 facts, with each fact associated with [1, 2, 3, 5, 10] paraphrased statements. We subsequently compared the retention rates of both edited and intrinsic knowledge following SFT on the Tulu-3-SFT-mixture dataset. The experimental results are summarized in Figure 5. Further details regarding model editing and fine-tuning hyperparameters are provided in Appendix B.
The editing success rate and the retention of edited knowledge both improve with an increasing number of paraphrases per fact. In contrast, the retention rate of intrinsic knowledge remains stable at approximately $80\%$. Notably, beyond a certain threshold of paraphrase, the retention of edit knowledge surpasses that of intrinsic knowledge. We posit that this occurs because, during pre-training, each piece of intrinsic knowledge fact is acquired from a limited number of paraphrases. When the number of paraphrases for an edited fact meets or exceeds the average number of knowledge variants encountered during pre-training, its retention resilience becomes comparable to, or even greater than, that of intrinsic knowledge. Our experiments indicate that providing approximately 3 paraphrases is sufficient for edit knowledge to match the retention rate of intrinsic knowledge.
5.2 Fine-Tuning with Frozen Selected Layers
@@
For Locate-then-Edit method, the layer that the edited knowledge being inserted into is where the knowledge located  Meng et al. (2023a); Geva et al. (2021). We hypothesize that preserving edited knowledge can be improved by avoiding fine-tuning of layers containing the target knowledge. To test this hypothesis, we evaluate two layer-specific fine-tuning strategies on GPT-2 XL:
@@
Finetune only later layers. We freeze the early layers and fine-tune the layers beyond a specified threshold layer. As GPT-2 XL contains 48 transformer layers, we set the finetuning layer threshold to be 10, 20, 30, 40 and compare their result. From the Figure 6(a), we find that larger layer threshold can improve the edited knowledge retention rate while having similar intrinsic knowledge retention rate. If only layers after 40 are finetuned, the edited knowledge reaches the same level of knowledge retention with intrinsic knowledge. In addition, finetuning after a layer threshold has similar intrinsic knowledge retention rate with full finetuning retention rate, showing that this method does not impair the intrinsic knowledge.
@@
Finetune only layers in a window. We freeze all layers except a window of layers. We focus on two methods: ROME and FT, because both of them incorporate knowledge by editing a single layer. We set the window size to be 5. To reserve a larger room of layers for the experiment, for both FT and ROME, we choose layer 10 instead of 6. From the Figure 6(b), we observe that fine-tuning achieves lowest edited knowledge retention rate when the window centered at layer 10-where the edited knowledge is located-while exhibiting higher edited knowledge retention rate for window does not include the edit layer, or even centered farther from this layer. In contrast, intrinsic knowledge maintains a similar knowledge retention rate across different window center positions.
6 Conclusion and Future Work
We examine how different downstream fine-tuning tasks affect previous edited knowledge. First, we demonstrate that knowledge incorporated via KE methods are particularly sensitive to downstream fine-tuning data, and none of ROME, MEMIT, FT and MALMEN can edit the knowledge to retain as intrinsic knowledge that suit for all the downstream fine-tuning tasks: they usually output related tokens that belong to the same category rather than the true target token.
Finally, we propose two solutions to improve the resilience of edited knowledge, the first approach increases the number of paraphrased expressions during the editing phase, which fundamentally make the edited knowledge comparable to intrinsic knowledge, the second implements a practical fine-tuning adaptation through selective layer freezing, preserving edits by avoiding modifications to layers most critical to the edited content.
For future KE methods, post-edit retention should be assessed not only immediately after editing, but also after fine-tuning. An interesting research direction involves developing editing techniques that automatically instill knowledge with a high retention rate, potentially leveraging models’ paraphrasing capabilities to create edits inherently resistant to fine-tuning drift.
@@ 7 Limitations
While our layer-freezing approach provides a solution for preserving single-edited knowledge, it introduces key limitations. Firstly, it restricts the model’s ability to acquire new knowledge during fine-tuning. Secondly, this method would decrease the training efficiency. Lastly, it cannot handle the extreme case that multiple edits that all layers become occupied by prior edits. As a result, development of a method for preserving multiple edits knowledge without compromising model plasticity or training efficiency would be an important future direction.
Appendix A Additional Related Work
Memory-based methods store edits in a explicit memory without modifying the model’s parameters  (Mitchell et al., 2022b; Zheng et al., 2023; Hartvigsen et al., 2023). For example, SERAC is a gradient-free memory-based model editing method that stores edits in an explicit memory and uses a scope classifier to determine if a test input is within the scope of any cached edits. If within scope, a counterfactual model predicts the label based on the most relevant edit example; otherwise, the base model’s prediction is used  (Mitchell et al., 2022b).
Appendix B Dataset Details
Table 3: Information for fine-tuning dataset.
@@ Appendix B Dataset Details
Appendix C Implementation Details
C.1 Knowledge Editing
To enhance locality, we fine-tune only a specific layer while freezing all others, as this approach demonstrates superior locality compared to full-model fine-tuning (Gangadhar and Stratos, 2024). A study has been conducted to determine the optimal layer for knowledge edits  (Hase et al., 2023; Meng et al., 2023a). For FT and ROME method, we choose layer 1 and layer 6 to be the edit layer respectively.
@@ C.1 Knowledge Editing
Full fine-tuning for editing(FT): We choose Adam optimizer with learning rate of 5e-5, maximum training step of 25, weight decay of 0 and early stopping loss of 0.01. We fine-tune one specific layer’s $mlp_{proj}$ of the model. We choose layer 1 as the edit layer for both GPT-2 XL and Llama3-8B.
@@
ROME: We emply the same hyper-parameters for ROME as the setting in original paper  Meng et al. (2023a): We choose learning rate of 0.5, maximum training step of 50, weight decay of 0.5 and KL factor of 0.0625. We perform model edition on one layer of the model. We choose layer 6 as the edit layer for both GPT-2 XL and Llama3-8B.
MEMIT: We choose learning rate of 0.2, maximum training step of 50, weight decay of 0.003, editing layers ranging from 3 to 8 and KL factor of 0.0625. We edit 10 facts in each model edition, so we perform 5 model edition experiments for each fine-tuning task.
@@
MALMEN: We adopt the same hyper-parameters for MALMEN as those used in the original paper  Tan et al. (2023). We observe that choosing later layers can achieve better edition success rate. We select the model editing hyperparameter such that the edited model can predict true target for all the prompts in our edit dataset. Specifically, for GPT-2 XL, we edit layers ranging from 43 to 48(out of 48), while for Llama3-8B, we edit layers ranging from 27 to 32(out of 32).
AlphaEdit: We adopt the same hyper-parameters for AlphaEdit as those used in the original paper  Fang et al. (2024). We select the model editing hyperparameter such that the edited model can predict true target for all the prompts in our edit dataset. For GPT-2 XL, we edit layers ranging from 13 to 17(out of 48).
C.2 Downstream Fine-tuning Tasks
We perform data filtering on the training set of all the downstream task, only keep the data that are irrelevant to the edit knowledge. To fairly compare these edit methods, it is essential to develop a metric to quantify the extent to which fine-tuning influences the model. To ensure that the fine-tuning impact is consistent across the four methods (FT, ROME, MEMIT and MALMEN), for each of these downstream fine-tuning tasks, we sample and fix an evaluation dataset from the fine-tuning dataset and standardize the training stopping criteria, and employ the same fine-tuning hyperparameter.
For a fine-tuning task, ensuring consistent fine-tuning impacts across different post-edit models is challenging but critical for fairness. We set the same stopping criteria for fine-tuning different post-edit models. Some experiments take more training epoch to reach the stopping criteria. For example, we note that fine-tuning with larger layer threshold needs larger training epoch. Notably, experiments with larger layer thresholds require more training epochs to meet these criteria, as fewer parameters are updated, demanding greater training effort to incorporate the same volume of knowledge into the model. For FT and ROME, we also try different edited layer. We find that for both of the methods, edit early layer (prior to layer 10) can achieve similar result.
Next token prediction on general text To systematically assess the impact on both the model’s intrinsic knowledge and the edited knowledge, we employed an unstructured dataset distinct from all pre-training corpora. Specifically, for our experiments on the GPT-2 XL, which was pre-trained on webtext, we choose the Common Crawl dataset for fine-tuning. Since a small subset of the data suffices to demonstrate the influence, we sampled 60k data for our training set. We evaluate the fine-tuning influence on a validation set, which contains of 1k instances sampled from the training set. We employ loss as a metric to measure the volume of knowledge acquired from fine-tuning. The fine-tuning is stopped once the validation loss goes below the threshold of 3.1.
Object prediction for factual knowledge triplets To investigate the influence of structured factual data on model behavior, we construct a fine-tuning dataset that mirrors the structure of our edit data while containing distinct factual content. Specifically, we sample data from COUNTERFACT train split, which has no overlap with the one used in upstream edit. We select only instances where the pre-trained model fails to predict the true target, and sample 3,000 of them. For supervision, we use the true target from each example as the training label. To monitor progress, we evaluate model performance on a validation set, which contains 100 instances sampled from the training set. Training terminates once the model achieves above 70% accuracy on this validation set.
Sentiment classification task To assess the impact of downstream fine-tuning on pre-trained knowledge, we employ a classification task as our benchmark evaluation. Specifically, we utilize the IMDB sentiment analysis dataset which consists of 25k movie reviews paired with binary sentiment labels. For the classification architecture, we append a fully connected layer to the final hidden state of the end-of-sequence (EOS) token. Classification accuracy on the validation set can be a fair metric, as it is the target index people are aiming to improve. The validation set contains 100 samples, which has no overlap with the training set. We fine-tune the model until the model achieves a classification accuracy exceeding 95% on the validation set.
Instruction fine-tuning task Supervised fine-tuning has emerged as a prevalent downstream adaptation method for LLMs. In this paradigm, each training instance consists of an instruction (question) paired with its corresponding response (answer). For our experiments, we employ the Tulu-3-SFT-mixture dataset  Lambert et al. (2024) for instruction fine-tuning, which consists of 10k movie reviews paired with binary sentiment labels. To assess the training influence, we evaluate model performance on a validation set, which contains 1k instances sampled from the training set. Training terminates once the model achieves above 70% accuracy on this validation set. We evaluate the training progress by compute the model’s loss on the validation set. We set the stopping criteria as validation loss goes below the threshold of 2.5.
C.3 Hyperparameter for fine-tuning